OpenStudy (anonymous):

Which of the following could be the first step in loving the equation 4*=17

OpenStudy (anonymous):

4x=17, or 4^x=17?

OpenStudy (anonymous):

4^x=17

OpenStudy (anonymous):

do you know logs?

OpenStudy (anonymous):

no not really

OpenStudy (anonymous):

4^x = 17 so: log(17)/log(4) = x We know that 4^2 = 16, so x is 2.something. log(17)/log(4) = 2.043731420625169704127033005405202177005633641172410344063304...

OpenStudy (anonymous):

We would just write x = log(17)/log(4)

OpenStudy (jdoe0001):

"Which of the following could be the first step in \(\bf loving\) the equation 4*=17 " well, look it... is it lovely?

OpenStudy (jdoe0001):

so, that'd be the first step =)